Problem 7-13
Nonconstant Growth Stock Valuation

Simpkins Corporation does not pay any dividends because it is expanding rapidly and needs to retain all of its earnings. However, investors expect Simpkins to begin paying dividends, with the first dividend of $0.75 coming 3 years from today. The dividend should grow rapidly - at a rate of 80% per year - during Years 4 and 5. After Year 5, the company should grow at a constant rate of 4% per year. If the required return on the stock is 15%, what is the value of the stock today (assume the market is in equilibrium with the required return equal to the expected return)? Round your answer to the nearest cent. Do not round your intermediate computations.

$  

Problem 7-12
Nonconstant Growth Stock Valuation

Assume that the average firm in your company's industry is expected to grow at a constant rate of 4% and that its dividend yield is 5%. Your company is about as risky as the average firm in the industry and just paid a dividend (D0) of $1.5. You expect that the growth rate of dividends will be 50% during the first year (g0,1 = 50%) and and 25% during the second year (g1,2 = 25%). What is the required rate of return on your company’s stock? What is the estimated value per share of your firm’s stock?

$  
Answer #1

1.
=0.75/1.15^3+0.75/1.15^3*(1.8/1.15)+0.75/1.15^3*(1.8/1.15)^2+0.75/1.15^3*(1.8/1.15)^2*(1.04/(15%-4%))=13.89555303

2.
required return=4%+5%=9%
Value per share=1.5*(1.5/1.09)+1.5*(1.5/1.09)*(1.25/1.09)+1.5*(1.5/1.09)*(1.25/1.09)*1.04/(9%-4%)=53.66972477